The ssqquote Package for L^{A}TEX 2ε

Copyright (C) 1994 by Ulrik Vieth

November 28, 1995

1 Introduction

This contributed package for L^{A}TEX 2ε provides the necessary font declarations needed to access the cmssq font family, i.e. Computer Modern Sans Serif Quotation Style, in terms of NFSS.^{1} It also provides a little example package file that shows how to define some appropriate font changing commands.

\ssqfamily Once you have installed the font definition files OT1cmssq.fd and T1cmssq.fd

\textssq provided here, you can use low-level NFSS commands to access the cmssq font family in your documents, even if you do not plan to use the example package ssqquote.sty. Apart from defining the font changing commands \ssqfamily and \textssq, that package file also provides an example application that uses the cmssq font family.

chapterquotes The chapterquotes environment provided as an example is based on ideas used in the \endchapter macro of the manmac format that was used by DEK in the TEX and manuals. Its purpose is to typeset a few nice quotations at the end of chapters using the cmssq font family in a smaller typesize than the regular text. While keeping the distinctive ragged-left formatting from the \endchapter macro, we have tried to make the chapterquotes environment more flexible, optionally allowing quotations to be placed at the top of the next page if there isn’t enough room left at the bottom of the current page.

1Throughout this document, the term NFSS should be understood to refer to release 2 of NFSS, i.e. the release that’s part of L^{A}TEX 2ε and that’s intended to become part of L^{A}TEX3 in the future.